Find the measure of an angle if its supplement measures 10 degrees more then three times its complement.
Answer by flame8855(424)

let m is the angle s be the supplement and c be the compliment
s=10+3c
180-m=10+3(90-m)
180-m = 10 +270 -3m
-100 = -2m
m = 50